CRY+VIS Interface Meeting
Date: 2021/6/11 13:00-13:50
Participants: Takahashi, Ikeda, Miyo, Sato, Ushiba, Tomaru, Hirata, Aso, Akutsu, Yokozawa, Kimura
- IX: H4 geophone should be replaced. For alignment of IFO, it is necessary to enter the cryostat.
- IY: Cryostat was closed on 6/7. Cryo-payload was released on 6/10.
- EX: Adjustment of Yaw FR was started, but not yet finished. Repair of cable is necessary. Closing top chamber by 6/17. Green beam will be introduced.
- EY: F0 and IP were tuned. Cleaning with pure water was tested.
- Fizeau interferometer could not been used.
- The measurement of cleaned viewport is under preparation.
- Smear was found with strong LED light. It went to the bath again.
- Making the handling jig takes more time, maybe until the next-next week.
- Corrugate tube fixing jig: Other 3 sets will be ready by the end of June.
- TM height setting: Next target is ITMY.
- Commissioning in air: Damping of PR3 was implemented. Auto TF measurement was made.
- Commissioning in vacuum:
- Overall commissioning:
- Acceptance check: Group work is set every Friday. Maintenance day was also set on Friday.
- PF actuator: After SC.
- Jig for FC painting:
- Separation of rack in 2F: On the maintenance days.
- Disassembly of PRM/PR2 assembly frame: This will be asked to MPC in early July. It is reasonable to apply at the same time as the fixing jig.
Participants: Takahashi, Fabian, Hirata, Miyo, Sato, Aso、Terrence, Ishizaki, Sugiyama, Tanaka
This is generated by python script in /users/Miyo/dropbox/Script/scrapping/main.py.
- Tuned the IP and F0 in EY.
- Adjusting the F0 FR in EX.
- Released the payload and adjusted the filter chain in IY.
- Assembled the additional geophone pods.
We lowered the SR pre-isolators: when the GAS filters areat zero, the SR2, SR3 and SRM mirrors are at approximately 50, 100 and 300 um above their respective correspoding nominal heights. The laser level thickness is around 2 mm, so the uncertainties are larger than the values of the measured heights.
- We also adjusted the IP-L and IP-T finely with the jacks after lowering the pre-isolators.
- In the three SR suspensions
- We set the F0 FRs hardware limit switches in suitable positions,
- Adjusted the amount of ballast mass on the F0 keystone taking into account the buoyancy corrections and
- Installed and tested the thermistors.
- In the BS, Hirata-san and Washimi-san checked the in-vacuum cable.
- Additionally, following Ikeda-san's instructions, we released the SRM F1 FR, which was jammed.
What's left to do?
- Tune SR IPs resonant frequencies (3 days).
- Test the stability of the BS IP by checking it comes back after large excursions (1 day or less).
- Fix SR2 and SR3 IP dysfunctional limit switches (1 day or less).
- Check BS height for the sake of information (1 day or less).
We aim to finish next week.
- Type-A team: IY(F0 tuning), EX(Cable repair, Yaw FR adjustment, IP adjustment)
- Type-B team: SR(IP tuning)
IMC suspensions are ready in air  will finish 2021/04/21.
PRs are ready in air  will finish 2021/07/09.
ETMX is ready in air  will finish 2021/08/25.
ITMX is ready in air  will finish 2021/08/31.
ETMY is ready in air  will finish 2021/09/16.
ITMY is ready in air  will finish 2021/11/08.
BS is ready in air  will finish 2021/11/24.
SRs are ready in air  will finish 2021/12/20.
Output suspensions are ready in air  will finish 2021/12/21.
TMSY is ready in air  will finish 2021/??/??.
TMSX is ready in air  will finish 2021/??/??.
The week from 6/14
- Takahashi: 6/14-18
- Sato: 6/14-18
- Hirata: 6/14-18
- Yano: 6/14-18
The week from 6/21
- Takahashi: 6/21-22
- Hirata: 6/21-22
- Yano: 6/21-25